OpenStudy (anonymous):

Determine the following relations for the van der Waals equation of state for real gases P=[(RT)/(V-b)]-[a/(V^2)], where a and b are constants that are experimentally fit to each gas. Determine alpha. Determine kappa.

OpenStudy (aaronq):

For what gases? is there more data?

OpenStudy (anonymous):

Nope this is the only information provided. Its asking for the equation that is determined not a numerical answer.

OpenStudy (aaronq):

what is alpha and kappa, though? no definition of them was provided

OpenStudy (anonymous):

Alpha was actually supposed to be Beta. I looked them up in the textbook and they are: \[\beta=\frac{ 1 }{ V }\left( \frac{ ∂V }{ ∂T} \right)_{P} \] and \[\kappa=-\frac{ 1 }{ V }\left( \frac{ ∂V }{∂P} \right)_{T}\]